OpenStudy (anonymous):

Will give medal! The altitude to the hypotenuse of a right triangle divides the hypotenuse into segments of lengths 6 and 9. What is the length of the altitude? A. 9 sqrt (2) B. 6 sqrt (6) C. 3 sqrt (6) D. 6 sqrt (3)

OpenStudy (michele_laino):

I think that:  from theorem of Euclide, we have: x^2=6*15, y^=9*15 now we can write: x*y=k*15=2*area of triangle, squaring both sides of the above equality, we can write: \[x ^{2}y ^{2}=k ^{2}*15^{2}\] please now continue

OpenStudy (anonymous):

Umm..I'm afraid I don't quite understand.

OpenStudy (michele_laino):

please substitute in the above equation, x^2, y^2

OpenStudy (campbell_st):

start with a diagram  you know a^2 + b^2 = 15^2 then if you look at the 2 smaller right triangles you will see 9^2 + d^2 = a^2 and 6^2 + d^2 = b^2 so subsituting you get 9^2 + d^2 + 6^2 + d ^2 = 15^2 now solve for d

OpenStudy (michele_laino):

using a theorem of Euclide we can write: x^2=6*15, and analogously: y^2=9*15 do you agree?

OpenStudy (campbell_st):

or 81 + 36 + 2d^2 = 225 117 + 2d^2 = 225 solve for d hope it helps

OpenStudy (anonymous):

Sorry, this is so confusing! I'm really bad at this stuff...

OpenStudy (campbell_st):

ok... so do you know pythagoras' theorem...?

OpenStudy (anonymous):

No, I don't remember learning that

OpenStudy (campbell_st):

wow... that makes it extremely hard

OpenStudy (anonymous):

yea...sorry

OpenStudy (campbell_st):

I'll leave my method there... so perhaps you can go and research pythagoras theorem and then comeback and study the solution to see if it makes sense

OpenStudy (michele_laino):

area of our right triangle is: A=(x*y)/2, or A=(hypotenuse*altitude)/2 so we can write: \[\frac{ x*y }{ 2 }=\frac{ hypotenuse*altitude }{ 2 }\] and then: \[x*y=hypotenuse*altitude\] now I called k our altitude, so substituting, we have: \[x*y=k*hypotenuse\] squaring both sides: \[x ^{2}y ^{2}=k ^{2}*(hypotenuse)^{2}\] since hypotenuse = 15: \[x ^{2}*y ^{2}=k ^{2}*15^{2}\] now inserting x^2 and y^2, we have: \[(6*15)*(9*15)=k ^{2}*15^{2}\] now, please, solve for k

OpenStudy (anonymous):

Okay, I will try that then.

OpenStudy (campbell_st):

Pythagoras's theorem is a Year 7 or 8 topic and is fundamental well working with right triangles.